OpenStudy (anonymous):

A chemist wishes to mix some pure acid with some water to produce 16 L of a solution that is 30% acid. How much acid and how much water should be mixed. can you write out the steps? cause I don't just wanna know the answer I wanna learn so I can do other problems like this. thanks :)

OpenStudy (campbell_st):

100% = 16L 30% acid = 0.3 x 16 70 water = 0.7 x 16 to check your answer add the two quantities together and you should get 16L

OpenStudy (anonymous):

okay sorry haha i'm confused easily, pure acid: 100%, then 16L with 30% is the solution, then you'll get two answers?

OpenStudy (mertsj):

Let x = amount of pure acid to add 100 % x = amount of acid in the pure acid 16-x= the amount of water to add 0%(16-x)= amount of acid in the water 30%(16)= amount of acid in the final solution 1.00x+0(16-x)=.30(16) x=4.8 L acid to add to 11.2 L of water

OpenStudy (anonymous):

OMG thank you much!!! people are so patient here I LOVE this website!